OpenStudy (anonymous):

If tan theta =12/5 and cos theta =-5/13, then what is sin theta?

OpenStudy (anonymous):

@peachpi

OpenStudy (michele_laino):

hint: \[\tan \theta = \frac{{\sin \theta }}{{\cos \theta }}\]

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

@Michele_Laino totally missed that. Much simpler than what I was going to do

OpenStudy (anonymous):

how do i go on from there

OpenStudy (anonymous):

what @Michele_Laino has posted is an identity for tan. From here you plug in your numbers and solve for sin Θ

OpenStudy (michele_laino):

ok! please continue @peachpi

OpenStudy (anonymous):

the answer i got is 12/13

OpenStudy (anonymous):

im not sure what i did, is that correct

OpenStudy (anonymous):

you should have gotten a negative number

OpenStudy (usukidoll):

you're missing a negative sign.

OpenStudy (anonymous):

so the answer would be -12/13???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just did the question again and got -5/13..... i am overthinking the problem i think

OpenStudy (anonymous):

so the answer is -12/13

OpenStudy (anonymous):